On January 1, 2017, Marin Company issued 10-year, $2,150,000 face value, 6% bonds, at par. Each $1,000 bond is convertible into 16 shares of Marin common stock. Marin’s net income in 2017 was $305,000, and its tax rate was 40%. The company had 97,000 shares of common stock outstanding throughout 2017. None of the bonds were converted in 2017.
Answer #1

Basic earning per share = Net income/Share outstanding

= 305000/97000

Basic earning per share = 3.14 per share

Diluted earning per share = Adjusted net income/Adjusted diluted share

Adjusted diluted shares = 97000+(2150000*16/1000) = 131400 shares

Adjusted net income = 305000+(2150000*6%*60%) = 382400

Diluted earning per share = 382400/131400 = 2.91 per shares